# Reinforcement Learning Teachers of Test Time Scaling

### v2

@@ D.1 RLT reward and R1 traces analysis
Appendix E Limitations and future implications
E.1 Limitations and unexplored directions
This work’s purpose was to introduce a new class of Reinforcement-Learned Teachers designed to avoid the exploration challenge of sparse rewards and align the optimization of RL-trained LMs with the test-time goal of downstream distillation. However, there are still several limitations and improvements which we hope will be tackled in future extensions. First, to output explanations, the RLT framework relies on access to the ground-truth solutions. Hence, when used with datasets and domains where this information might not be available or not be practical to recover without querying LMs, small RLTs might have to again rely on larger models, even though still to a lesser extent than prior distillation pipelines. Similarly, as described in Section 4.1 and Appendix A, our current training recipe does not exclusively involve RL, with an initial phase to familiarize our small model with its new teaching format and role, again relying on some level of initial access to pre-collected reasoning examples that can be used accordingly. We also note that RLT training makes use of an additional student model for computing the rewards. In practice, however, we note the downsides from this were minimal, as simple parameter offloading removed any potential memory burden of having this model on GPU memory during backpropagation, and actual training time was dominated ($>90\%$) by the costs of long-context autoregressive generation with our computational setup.

### v3

@@ D.1 RLT reward and R1 traces analysis
Appendix E Extended discussion and limitations
E.1 Traditional RL and the importance of overcoming reward sparsity
In traditional RL for robotics, neural network policies can learn to solve tasks from random initializations thanks to being guided by dense reward functions designed and tuned to provide domain-specific guidance on the level of progress [53, 54]. Without this guidance, they would be faced with an infeasible exploration challenge made exponentially more difficult by the task horizon [55]. This is because dense rewards allow the policy gradient optimization to rank the relative progress obtained across its initial suboptimal actions, allowing the policy to bootstrap from partial solutions and extrapolate far beyond its initial capabilities. However, in the context of traditional LM reasoning, as correctness-based rewards are inherently sparse, this extrapolation is not possible. In particular, if an LM is too weak to provide the optimal solution for a task, all its rewards (and, thus, the policy gradient) will be zero. In contrast, we note this limitation would not apply to our RLT, which can make use of a dense reward function and obtain a learning signal to iteratively improve even when all sampled initial explanations are suboptimal, as with the traditional RL framework.
E.2 Limitations and unexplored directions
This work’s purpose was to introduce a new class of Reinforcement-Learned Teachers designed to avoid the exploration challenge of sparse rewards and align the optimization of RL-trained LMs with the test-time goal of downstream distillation. However, there are still several limitations and improvements which we hope will be tackled in future extensions. First, to output explanations, the RLT framework relies on access to the ground-truth solutions. Hence, when used with datasets and domains where this information might not be available or not be practical to recover without querying LMs, small RLTs might have to again rely on larger models, even though still to a lesser extent than prior distillation pipelines. Similarly, as described in Section 4.1 and Appendix A, our current training recipe does not exclusively involve RL, with an initial phase to familiarize our small model with its new teaching format and role, again relying on some level of initial access to pre-collected reasoning examples that can be used accordingly. We also note that RLT training makes use of an additional student model for computing the rewards. In practice, however, we note the downsides from this were minimal, as simple parameter offloading removed any potential memory burden of having this model on GPU memory during backpropagation, and actual training time was dominated ($90\%$) by the costs of long-context autoregressive generation with our computational setup. Recent work has shown early promise in overcoming this inherent bottleneck of RL methods by grounding the model’s reasoning process away from self-generated tokens, and using separate spaces such as diffusion [56]. Yet, while this complementary approach can be more efficient, it has yet to have the same impact as RL reasoning at larger scales.